# The Nontrivial Vacuum Structure of an Extended $t\overline{t}$ BEH (Higgs) Bound state

### v3

@@ IV Full Action
The full action in the pointlike limit then becomes that of the standard model BEH boson coupled to top quarks with in the unprimed parameters:
@@
IV.1 BEH Potential
Note the theory generates the usual “sombrero potential”:
@@ IV.1 BEH Potential
$H^{\prime}(X){\Phi}(r)=H(X^{\mu})\widetilde{\Phi}(r)\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}\widetilde{\Phi}(r^{\mu})+\frac{h(X,r^{\mu})}{\sqrt{2}}\\ 0\end{array}\biggr),$ (57)
where $h(X,r^{\mu})=h(X)\widetilde{\Phi}(r^{\mu})$ is the physical BEH boson and $\pi^{a}(X)$ are Nambu-Goldstone bosons, which are phase factors of $H(X^{\mu})$. The electroweak symmetry is broken spontaneously and the gauge fields absorb the Nambu-Goldstone phase factors and acquire mass in the usual way. The resulting broken phase of the BEH field, $h(X)$ is then the standard model and the canonical normalization of $h(X)$ follows from the $Z$ normalization of the internal field $\widetilde{\Phi}(r)$ in eq.(27).
The Nambu-Goldstone modes have only $X$ dependence, and the neutral component of $H^{\prime}$ is then
$H^{\prime}(X)\Phi=H(X^{\mu})\widetilde{\Phi}(r)\rightarrow v_{weak}\widetilde{\Phi}(r)+h(X)\widetilde{\Phi}(r)/\sqrt{2},$ (58)
where $h(X)$ is the SM BEH boson. The “two-body” field, $h(X)$, is now associated with apparent relative time through $\widetilde{\Phi}(r)$, however, the key feature is that $\widetilde{\Phi}(r)\sim\widetilde{\Phi}(0)+{\cal{O}}(r_{\mu}r^{\mu})+$... is a Lorentz invariant function of $r_{\mu}r^{\mu}$ and makes no reference to a particular $\omega^{\mu}$. This expansion can, in principle, yield a form factor in BEH interactions and/or higher dimension operators as described above. Hence the SM BEH field, $h(X)$ is a collective object and its precise two-body nature is actually blended with the collective vacuum $\Phi$. Likewise, the Nambu-Goldstone bosons are “eaten” by the gauge fields in the usual way, and $\Phi(r^{\mu})$ is integrated out and does not affect the usual combinatorial.
The presence of the collective wave-function $\Phi(r^{\mu})$ is not detectable in the kinetic terms, nor in the interaction in the pointlike limit. We obtain the BEH kinetic term eq.III.1) of the SM :
$+\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}\rightarrow\frac{1}{2}ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;(\partial h(X))^{2}|\widetilde{\Phi}(r)|^{2}=\frac{1}{2}\int\!\!d^{4}X(\partial h)^{2},$ (59)
where we integrate out $\widetilde{\Phi}$ in eqs.(III.1,59) using eq.(40). In fact, there is no way to discern the compositeness of the BEH field from the kinetic and mass terms (this requires the Yukawa interaction and quartic terms). Moreover, $\langle H\rangle\rightarrow v_{weak}$ with the covariant derivative, $D_{\mu}$, of eq.(80), leads to:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z|DH(X^{\mu})|^{2}|\widetilde{\Phi}(r)|^{2}\biggr)\rightarrow$
$=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(M_{W}^{2}W^{+}W^{-}+\frac{1}{2}M_{Z}Z^{2}\biggr)|\widetilde{\Phi}(r)|^{2}=\!\int\!\!d^{4}X\biggl(M_{W}^{2}W^{+}W^{-}+\frac{1}{2}M_{Z}Z^{2}\biggr).$ (60)
Hence we generate the $W^{\pm}$ and $Z^{0}$ mass terms in the usual way, and the Nambu-Goldstone bosons have become their longitudinal gauge components.
The usual effective action for the BEH boson $h(X)$ emerges:
@@
The structure of the collective field, $\Phi(r^{\mu})$, can in principle be probed through the extended interactions, but the effects will be suppressed. The Yukawa coupling term in natural top condensation takes the form chris1 ; chris2 in the broken phase:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=\sqrt{2N_{c}J}g^{2}_{0}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r^{\mu})\;\biggl(v_{weak}+\frac{h(X)}{\sqrt{2}}\biggr)\widetilde{\Phi}(r^{\mu}){+h.c.},$ (62)
where $\psi_{L}=(t,b)_{L}$ and $\psi_{R}=t_{R}$ (and $J=16$ is the Jacobian in passing from coordinates $(x,y)$ to $(X,r)$ chris1 ).
To a good approximation we can take the pointlike limit of the potential, and we have $D_{F}(2r^{\mu})\rightarrow(JM_{0}^{2})^{-1}\delta^{4}(r)$. Hence
$g_{Y}=\hat{g}_{Y}|\widetilde{\Phi}(0)|,\qquad\makebox{where,}\qquad\hat{g}_{Y}\equiv g_{0}^{2}\sqrt{2N_{c}/J}=g_{0}^{2}\sqrt{3/8}.$ (63)
This then leads to the conventional mass term for the top quark and BEH coupling:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=g_{Y}\int\!\!d^{4}X\;[\overline{\psi}_{L}(X)\psi_{R}(X]_{f}\biggl(v_{weak}+\frac{h(X)}{\sqrt{2}}\biggr){+h.c.}.$ (64)
@@ IV.2 Integrating Out $r^{\mu}$ and Higher Dimension Operators
If we go beyond the pointlike limit in the BEH-Yukawa interaction we will have Lorentz invariant $O(1/M_{0}^{2})$ corrections. Before taking the pointlike limit of the Yukawa interaction, as in eq.83), we can rewrite this in the broken phase as:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=m_{top}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}(M_{0}^{2}D_{F}(2r^{\mu}))\;F(r^{\mu})+h.c.,\qquad F=\frac{\widetilde{\Phi}(r)}{\widetilde{\Phi}(0)}\sim 1+a_{1}r^{2}+$... (65)
Note that,
$\!\!\!\!\!\!\!\!\!\!(D_{F}(2r^{\mu}))(r^{p})\sim\int\!\!\frac{d^{4}q}{(2\pi)^{4}}\frac{((\partial_{q})^{p}e^{iqr})}{q^{2}-M_{0}^{2}}\sim\frac{1}{M_{0}^{2+p}}\delta^{4}(r),$ (66)
implying that the expansion in $r$ corresponds to an expansion in $M_{0}^{-1}$, and derivatives $\partial_{X}^{p}$. The expansion in $r^{\mu}$ in all terms in the integrand generates a tower of assorted Lorentz invariant higher dimension operators such as:
@@
$\overline{t}\rightarrow b+W+(g,\gamma,Z))\qquad\text{and,}\qquad\overline{t}\rightarrow t+(g,\gamma,Z)).$ (68)
@@
$(\ell^{+}\ell^{-})\rightarrow(\gamma^{*},Z^{*})\rightarrow t+b+(W,g,\gamma,Z)).$ (69)
@@ V Summary
We have proposed a nontrivial vacuum for the natural top condensation theory. The vacuum is manifestly Lorentz invariant, composed of a collective Lorentz invariant sum over internal wave-functions, $\Phi=(1/\sqrt{N})\sum_{\omega}\phi_{\omega}(r^{\mu})$. Though each internal wave-function, $\phi_{\omega}$, is independent of relative time, $\tau$, they each have dependence upon the arrow of relative time, $\omega^{\mu}$, and are solutions to the Schrödinger-Klein-Gordon equation with eigenvalue $|\mu^{2}|$. The sum over $\phi_{\omega}$ makes the collective field, $\Phi(r^{\mu})$, Lorenz invariant and completely independent of $\omega^{\mu}$.
@@
Much can be done to explore this theory further, including the bilocal formalism itself. For a more complete theory, we note that the $b_{R}$ quark also participates in topcolor in an anomaly-free scheme, e.g., see Topcolor , and this should be adapted to the present dynamics. An additional $Z^{0}{{}^{\prime}}$ interaction can be introduced that makes the $\overline{b}b$ channel sub-critical, hence non-binding. Light fermion masses are presumably then generated in analogy to extended technicolor models Technicolor . So far, we have relied upon the intuition of the 1990’s topcolor scheme, but a more complete natural top condensation theory, including the light particle masses and mixing angles, could be readily formulated.
The theory is therefore testable, mainly by direct discovery of the octet of colorons at $M_{0}\approx 6$ TeV. The theory points toward an $SU(3)\times SU(3)\times SU(2)\times U(1)_{Y}$ gauge structure emerging at the $\sim 10$ TeV scale with likely additional $U(1)^{\prime}s$. The theory also offers sensitive probes of new contact interactions involving $t$- and $b$-quarks, and though flavor mixing there may be induced rare processes involving the other quarks and leptons Technicolor .
@@
Appendix A Orthogonality of $\phi_{\omega}(r)$
@@
We consider small $\theta$, hence $\phi_{\omega^{\prime}}(r^{\mu})\approx\phi(r^{0}\theta^{2},\vec{r})$. The overlap integral is dominated by the large $r=|\vec{r}|$ component and in this limit with eq.(75), and flat direction $r^{0}$,

### v4

@@ IV Full Action
IV.1 Recovery of Standard Model in Pointlike Limit of the Potential
The full action in the pointlike limit of the potential, $D_{F}(2r)\rightarrow\delta^{4}(r)/16M_{0}^{2}$, then becomes that of the standard model BEH boson coupled to third generation (top) quarks with the unprimed parameters. While $\Phi(r^{\mu})$ remains extended, it is integrated out, by its normalization, and relative time does not appear:
@@ IV.1 Recovery of Standard Model in Pointlike Limit of the Potential
The theory generates the usual “sombrero potential”:
@@
$H^{\prime}(X^{\mu}){\Phi}(r)=H(X)\widetilde{\Phi}(r)\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}+\frac{h(X)}{\sqrt{2}}\\ 0\end{array}\biggr)\widetilde{\Phi}(r),$ (60)
where $h(X)$ is the physical BEH boson and $\pi^{a}(X)$ are Nambu-Goldstone bosons, which are phase factors of $H(X^{\mu})$. The electroweak symmetry is broken spontaneously and the gauge fields absorb the Nambu-Goldstone phase factors and acquire mass in the usual way. The resulting broken phase of the BEH field, $h(X)$, is then as in the standard model and the canonical normalization of $h(X)$ follows from the $Z$ normalization of the internal field ${\Phi}(r)$ in eq.(33). Hence the SM BEH field, $h(X)$, is a collective object and its precise two-body nature is actually blended with the collective vacuum $\Phi$. Likewise, the Nambu-Goldstone bosons are “eaten” by the gauge fields in the usual way.
The presence of the collective wave-function $\Phi(r^{\mu})$ is not detectable in the kinetic terms, nor in the interaction in the pointlike limit of the potential. We obtain the BEH kinetic term from eq.(III.1) as in the SM:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}\rightarrow\frac{1}{2}ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;(\partial h(X))^{2}|{\Phi}(r)|^{2}=\frac{1}{2}\int\!\!d^{4}X(\partial h)^{2},$ (61)
where we integrate out $\Phi$} using eq.(52). In fact, there is no way to discern the compositeness of the BEH field from the kinetic and mass terms (this requires the Yukawa interaction and quartic terms below). Moreover, $\langle H\rangle\rightarrow v_{weak}$ with the covariant derivative, $D_{\mu}$, of eq.(81), and eq.(52) leads to:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z|DH(X^{\mu})|^{2}|{\Phi}(r)|^{2}\biggr)\rightarrow$
$=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(M_{W}^{2}W^{+}W^{-}+\frac{1}{2}M_{Z}Z^{2}\biggr)|{\Phi}(r)|^{2}=\!\int\!\!d^{4}X\biggl(M_{W}^{2}W^{+}W^{-}+\frac{1}{2}M_{Z}Z^{2}\biggr).$ (62)
Hence we generate the $W^{\pm}$ and $Z^{0}$ mass terms in the usual way, where the Nambu-Goldstone bosons have become their longitudinal gauge components. The usual effective action for the BEH boson $h(X)$ emerges:
@@ IV.2 Integrating Out $r^{\mu}$ and Higher Dimension Operators
The structure of the collective field, $\Phi(r^{\mu})$, can in principle be probed through the extended Yukawa and quartic interactions, but the effects will be suppressed by powers of $M_{0}$. The Yukawa coupling term in natural top condensation takes the form chris1 ; chris2 in the broken phase of eq.(III.1) (recall there is no $Z^{\prime}$ factor on the interactions and $H^{\prime}\Phi=H\widetilde{\Phi}$):
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=\widehat{g}_{Y}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r^{\mu})\;\biggl(v_{weak}+\frac{h(X)}{\sqrt{2}}\biggr)\widetilde{\Phi}(r^{\mu}){+h.c.},$ (64)
In the pointlike approximation of the potential and we have $D_{F}(2r^{\mu})\rightarrow(16M_{0}^{2})^{-1}\delta^{4}(r)$, then $\widehat{g}_{Y}\widetilde{\Phi}(0)=\widehat{g}_{Y}\phi_{\omega}(0)=g_{Y}$, which then leads to the conventional mass term for the top quark and BEH coupling:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=g_{Y}\int\!\!d^{4}X\;[\overline{\psi}_{L}(X)\psi_{R}(X]_{f}\biggl(v_{weak}+\frac{h(X)}{\sqrt{2}}\biggr){+h.c.}\qquad\text{where,}\qquad m_{top}=g_{Y}v_{weak}.$ (65)
If we go beyond the pointlike limit we can integrate out $r^{\mu}$ and the BEH-Yukawa interaction will generate $O(1/M_{0}^{2p})$ corrections. Note that if only a single component field, $\phi_{\omega}(r)$, comprised the entire wave function, e.g., $\Phi(r)=\phi_{\omega}(r)$, then we would generate Lorentz violating higher dimension operators at this stage. However, with the collective field $\Phi(r)$ that sums over all $\phi_{\omega_{i}}(r)$, as in eq.(4), these effects are now bona fide Lorentz invariant.
Rather than taking the pointlike potential limit, we can rewrite eq.66) in the broken phase as:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=m_{top}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}(M_{0}^{2}D_{F}(2r^{\mu}))\;G(r^{\mu})+h.c.,\qquad G(r^{\mu})=\frac{\widetilde{\Phi}(r)}{\widetilde{\Phi}(0)}\sim 1+a_{1}r^{2}+$... (66)
Likewise, we can expand $\overline{\psi}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]\sim\sum_{p}c_{p}r^{p}\overline{\psi}_{L}(X)D_{X}^{p}\psi_{R}(X)$ where $D_{X}$ are covariant derivatives wrt $X$. Hence the integrand becomes a power series in $r^{p}$. Then, integrating by parts, we have:
$\!\!\!\!\!\!\!\!\!\!\int d^{4}r\;C_{p}(X)r^{p}D_{F}(2r^{\mu})\sim\int d^{4}r\int\!\!\frac{d^{4}q}{(2\pi)^{4}}\;C_{p}(X)e^{iqr}\frac{\partial^{p}}{\partial q^{p}}\biggl(\frac{1}{q^{2}-M_{0}^{2}}\biggr)\sim C_{p}(X)\frac{1}{M_{0}^{2+p}},$ (67)
where we take the pointlike limit in the last term. Hence, the expansion in $r$ corresponds to an expansion in $M_{0}^{-1}$. The expansion in $r^{\mu}$ in all terms in the integrand generates a tower of assorted Lorentz invariant higher dimension operators such as a leading term:
@@
$\overline{t}\rightarrow b+W+(g,\gamma,Z)\qquad\text{and,}\qquad\overline{t}\rightarrow t+(g,\gamma,Z).$ (69)
@@
$(\ell^{+}\ell^{-})\rightarrow(\gamma^{*},Z^{*})\rightarrow t+b+(W,g,\gamma,Z).$ (70)
@@ V Summary
We have proposed a nontrivial vacuum for the natural top condensation theory. The vacuum is manifestly Lorentz invariant, composed of a collective Lorentz invariant sum over internal wave-functions, $\Phi=(1/\sqrt{N})\sum_{\omega}\phi_{\omega}(r^{\mu})$. Though each internal wave-function, $\phi_{\omega}$, is independent of relative time, $\tau$, they each have dependence upon the arrow of relative time, $\omega^{\mu}$, and are solutions to the Schrödinger-Klein-Gordon equation with eigenvalue $|\mu^{2}|$. The sum over $\phi_{\omega}$ makes the collective field, $\Phi(r^{\mu})$, Lorentz invariant and independent of $\omega^{\mu}$. Residual dependence upon $r^{\mu}$ can then be integrated out leading to suppressed Lorentz invariant operator corrections.
@@
Much can be done to explore this theory further, including the bilocal formalism itself. For a more complete theory, we note that the $b_{R}$ quark also participates in topcolor in an anomaly-free scheme, e.g., see Topcolor , and this should be adopted into the present dynamics. An additional $Z^{0}{{}^{\prime}}$ interaction can be introduced that makes the $\overline{b}b$ channel sub-critical, hence non-binding.
The theory is therefore testable, mainly by direct discovery of the octet of colorons at $M_{0}\approx 6$ TeV. The theory points toward an $SU(3)^{\prime}\times SU(3)\times SU(2)\times U(1)_{Y}$ gauge structure emerging at the $\sim 10$ TeV scale with likely additional $U(1)^{\prime}s$. The theory also offers sensitive probes of new contact interactions involving $t$- and $b$-quarks, and though flavor mixing there may be induced rare processes involving the other quarks and leptons Technicolor .
The $\overline{b}_{L}b_{R}$ mass term can in principle arise from instantons in the $SU(3)^{\prime}$ sector as in Topcolor ; the third generation lepton $(\tau)$ mass may arise by extending $SU(3)^{\prime}\rightarrow SU(4)$, i.e., treating leptons as the fourth color. Light fermion masses are presumably then generated in analogy to extended technicolor models Technicolor . So far, we have relied upon the intuition of the 1990’s topcolor scheme, Topcolor , but a more complete natural top condensation theory, including the light particle masses can be formulated. Our theory implies, therefore, a novel gauge dynamics, compositeness, and possibly the entry point to a complete theory of flavor physics.
@@
Appendix A Orthogonality of $\phi_{\omega_{i}}(r)$
@@
We consider small $\theta$, hence in leading order in $\theta^{2}$ it is sufficient to use $\phi_{\omega^{\prime}}(r^{\mu})\approx\phi(r^{0}\theta^{2},\vec{r})$. The overlap integral is dominated by the large $r=|\vec{r}|$ component and in this limit with eq.(76), and flat direction $r^{0}$, using a cut-off $r^{0}=T$ on the integral: